# Out-of-distribution generalization via composition: a lens through induction heads in Transformers

### v1

@@ E.2 Details about PTHs and IHs in LLMs
E.3 Histograms for PTH/IH matching
In Figure 5(c)(d), we only presented the histogram plots based on GPT2. We show that results from other LLMs are consistent with our findings. In Figure 22 and 23, we present the histograms of diagonal scores and the histograms of subspace matching scores.
@@ E.3 Histograms for PTH/IH matching
E.4 Histograms for shuffling experiments
@@
E.5 Histograms for projection experiments
@@
Appendix F Additional Related Work
Large Language Models.
Large Language Models (LLMs) are typically based on the Transformer architecture and are characterized by their enormous number of parameters and extensive pretraining on vast datasets. Notable examples include LLaMA[80], ChatGPT [59], GPT4 [60] and Claude [8]. These models utilize various pretraining methods such as masked language modeling [20, 45], and autoregressive pretraining [16]. Researchers have investigated the effects of pretraining on language model performance. Adapting LLMs to various downstream tasks has garnered significant attention in the field. This adaptation can take many forms, including the use of adapters [36, 37], multitask fine-tuning [85, 92, 93], in-context learning [55, 21, 72], reinforcement learning from human feedback (RLHF) [61], and methods for accelerating inference [31, 90]. Each of these approaches aims to enhance LLM performance or efficiency for specific applications or domains, allowing these powerful models to be tailored to a wide range of tasks and requirements.
In-context Learning.
LLM exhibits a remarkable ability for in-context learning (ICL) [16], particularly for generative models. This ability allows models to construct new predictors for test examples based on a prompt containing a sequence of labeled examples, without requiring parameter updates. The behavior of ICL has been the subject of several empirical studies. [97, 34, 47] have formulated the problems and reported on its sensitivity. To enhance performance, [18, 54] have employed meta-training with an explicit in-context learning objective. On the theoretical front, [89, 28] have provided frameworks to explain the underlying mechanisms of in-context learning. Studies by [82, 3, 49, 94], using linear models, have demonstrated how transformers can represent gradient descent and perform linear regression. On contrastive to these works, we present an mechanistic analysis that illustrates how LLMs can demonstrate OOD capability in ICL.

### v2

@@ E.1 Details about PTHs and IHs in LLMs
E.2 Shuffle intervention experiment
Evaluation.
$p_{2L+t}\big{(}s_{t}^{\#}\big{|}({\bm{s}}^{\#},{\bm{s}}^{\#},{\bm{s}}_{<t}\big{)}.$ (10)
For GPT-2, we gather the above probability for $t\in\{6,7,8,\ldots,L\}$ and every input sequence. Finally, in the left plot of Figure 5(e) in our main paper, we compare the histogram of the prediction probabilities of the original model, the edited model with shuffling, and the edited random baseline.
E.3 Projection intervention experiment
Calculating bridge subspace.
As mentioned in our main paper, we calculated the bridge subspace of rank $r$ based on the top-$K$ ${\bm{W}}_{\mathrm{QK}}$ from the list of IHs. We chose to edit $25\%$ of all heads using the projection matrices ${\bm{V}}{\bm{V}}^{\top}$ or ${\bm{I}}_{d}-{\bm{V}}{\bm{V}}^{\top}$ because we believe that $25\%$ of all heads are likely to be relevant to the copying task. These heads are selected as the $25\%$ top-ranking heads according to the IH attention score.
Evaluation.
50 80 150 170 200 200 200 200 200 220
Table 3: The rank $r$ of ${\bm{V}}$ used in the projection calculation.
E.4 Histograms for PTH/IH matching
In Figure 7(c)(d), we only presented the histogram plots based on GPT2. We show that results from other LLMs are consistent with our findings. In Figure 23 and 24, we present the histograms of diagonal scores and the histograms of subspace matching scores.
@@ E.4 Histograms for PTH/IH matching
E.5 Histograms for shuffling experiments
@@
E.6 Histograms for projection experiments
@@
Appendix F Additional results for the Common bridge representation hypothesis
In Figure 7(e), 7(f), we only report the results of shuffle and projection intervention experiments on the Copying task. For the other three tasks, Fuzzy Copying, IOI, and ICL, we follow the measurement in D.2 and the intervention method in E.2, E.3. We show in in Figure 27 the results of experiments. The results are similar to those presented in Figure 7.
(a) Shuffle experiment: reduced probability on Fuzzy Copying task.
(b) Shuffle experiment: reduced probability on IOI task.
(c) Shuffle experiment: reduced probability on ICL task.
(d) Projection experiment: reduced accuracy on IOI task.
(e) Projection experiment: reduced accuracy on IOI task.
(f) Projection experiment: reduced accuracy on ICL task.
Figure 27: A set of shuffle experiments: reduced probability on the Fuzzy Copying task.
Appendix G Experiment details for the Scaling Experiment
Model details.
We list the models in the experiments.
  • Pythia-36M 6 layers, 8 attention heads, and a hidden size of 256.
  • Pythia-70M 6 layers, 8 attention heads, and a hidden size of 512.
  • Pythia-160M 12 layers, 12 attention heads, and a hidden size of 768.
  • Pythia-410M 24 layers, 16 attention heads, and a hidden size of 1024.
  • Pythia-1B 16 layers, 8 attention heads, and a hidden size of 2048.
  • Pythia-1.4B 24 layers, 16 attention heads, and a hidden size of 2048.
  • Pythia-2.8B 32 layers, 32 attention heads, and a hidden size of 2560.
  • Pythia-7B 32 layers, 32 attention heads, and a hidden size of 4096.